OpenStudy (anonymous):

Ace Formalwear buys tuxedos from a supplier and has an invoice amount of $27,700. The terms of the sale are 3/20, n/30. If Ace sends a partial payment of $7,400 on the discount date, what is the net amount still due?

OpenStudy (amistre64):

tell me the notation ... what does your 3/20 n/30 tell us about the conditions?

OpenStudy (amistre64):

got this out of google books ...

OpenStudy (anonymous):

I know that if you pay by the 20th of the month you will receive 3% cash discount and no discount after the 30th.

OpenStudy (amistre64):

thats a good interpretation of the notations ... now for the mathing. payment/(1-.03) = credit applied so whats left over is: total - credit

OpenStudy (amistre64):

so the 7400 paid in is worth 7400/.97 = 7628.87 there abouts

OpenStudy (anonymous):

I have another question for anyone who can help it will be appreciated.Diane's Pottery Shop orders china from a supplier for $240 including $30 freight charges. The terms of sale are 3/10, n/30 ROG. The china arrives on 11/21 and Diane wants to take advantage of the cash discount. a. By what date must the invoice be paid to take advantage of the discount?

OpenStudy (amistre64):

eom i know means end of month ,,, what does the rog tell us?

OpenStudy (anonymous):

It means the discount does not start until the buyer receives the product.

OpenStudy (amistre64):

the what is the date, 10 days after november 21st ?

OpenStudy (anonymous):

I wrote December 1 and it was correct but I am now having problems with the second half As the bookkeeper for Diane's how much will you send the supplier?

OpenStudy (amistre64):

97% of the total

OpenStudy (amistre64):

if discount is equal to net ... the remaining debt is 0 therefore: payment/.97 = full price payment = full price * .97

OpenStudy (anonymous):

It was incorrect but I really do appreciate you helping me with this problem.

OpenStudy (amistre64):

what did you use for the full price?

OpenStudy (anonymous):

Here is the answer to both 3/10, n/30 ROG means that buyers can take a 3% discount within 10 days after receipt of goods. Full amount is due between day 11 and day 30 if cash discount period is missed. With the receipt of goods (ROG) the cash discount period begins when buyer receives goods, not the invoice date. a. By what date must the invoice be paid to take advantage of the discount? Do not abbreviate the name of the month. 10 days after 11/21 that is 21th of November is 1 of December. To take advantage of the discount the invoice must be paid by 1st of December. b. As the bookkeeper for Diane's how much will you send the supplier? Freight must be subtracted from the gross before calculating a cash discount (240 - 30)(1 - 0.03) = 210*0.97 = $203.7